# A Prospect-Theoretic Policy Gradient Algorithm for Behavioral Alignment in Reinforcement Learning

### v2

@@ 2 $\uparrow$ ? $\uparrow$
[…]
@@
[…]
@@
[…]
@@
Implementation details. In both cases, the risk-neutral optimal solution (going around the city center) is also a local optimum for the risk-averse objective, and, because it is a shorter path, is easier to stumble upon by chance when exploring the MDP. This means we have to implement special measures to force exploration. The algorithm used as is is prone to get stuck from time to time in local minima on this example. It would seem that our $w$ function, which is aggressively risk-averse, hinders exploration. To mitigate this, we introduce an entropy regularization term that we add to the score function with a decaying regularization weight in the policy gradient found in Theorem 6, see appendix I for further details. We incorporate entropy regularization in the policy gradient as follows:
@@
I.6 Electricity Management
We consider an electricity management system for an individual home which has solar panels for producing electricity and a battery for storing energy. The intensity of the solar panel’s electricity production follows a sinusoidal function during daytime hours and vanishes at night. The home consumes a random and varying amount of electricity and can buy and sell electricity to the outer grid. The selling price varies during the day whereas the buying price is fixed and significantly higher than the selling prices. We use public data for selling prices recorded on a national electricity network. We consider a 24h time frame starting at 6 am and we divide the day into twelve two-hour time slots. For each time slot, the agent has to decide how much to buy or sell to the grid given the production, the battery’s charge, the price on the market and the consumption. We run the algorithm with three different objectives, changing the $w$ function: a risk-neutral one, a risk-averse one and a risk-seeking one (see Table 2). We consider a Gaussian policy in which the mean is parameterized by a neural network. We report the results for running our algorithm in Fig. (2) (right). The most rewarding time to sell our electricity is around 4pm (see electricity prices in appendix I.6, Fig. 22, right). However, selling too much too soon exposes us to the risk of falling short of battery during the night and risking to buy it later for a higher price.
The risk-averse policy avoids selling a lot of electricity and tends to keep it stored until the end of the day. Conversely, the risk-seeking policy aggressively sells energy when the markets are high at the cost of possibly having to buy it again later in the day. We can see on Fig. 2 (right), where we plot the distribution of total returns for various trained models with different $w$ functions and a few different random initializations each, that, as we would wish to see, the risk-averse policy has the distribution with the best left tail (worst cases are not too bad), the risk-seeking distribution has the best right tail (best cases are particularly good). The risk-neutral policy has the best mean value.
Figure 22: Electricity prices in a typical day, the blue line (right-hand side scale) records the electricity price on the European market, the shaded area (left-hand side scale) represents the total electricity production in France.
I.7 Trading in Financial Markets
@@
  • Rewards: The rewards we consider are given by the log values of the ratio of the portfolio valuations at times $t$ and $t-1\,.$ Borrowing interest rates and trading fees are also considered in the computation. The reward function can also be easily modified in the environment thanks to the implementation of the Gym Trading Environment which builds on the standard Gym environments.
@@
Results. We refer the reader to Fig. 23 and Fig. 24 below. We make a few observations:
@@
  • Influence of the parameter $\alpha$ in KT’s utility (Fig. 24): Observe that the exponent $\alpha$ in the utility distorts the function and shifts the returns significantly. Lower values of $\alpha$ lead to higher returns in this setting where the returns (as per the ratio definition of the reward) are smaller than 1. This parameter $\alpha$ provides a degree of freedom to model the behavior of the agent as per their perception of the returns. Different values of $\alpha$ modify the curvature of the utility function (w.r.t. the reference point which is $x_{0}=0$ here) which is concave for gains and convex for losses.
Figure 23: Performance of our PG algorithm on a financial trading application. KT refers to Kahneman and Tversky’s utility function, x0 is the reference point used in that utility, exp. refers to exponential. Shaded areas are interquantile (25-75%) margins and curves report the median values over 10 different runs.
Figure 24: Performance of our PG algorithm on the same financial trading application. KT refers to Kahneman and Tversky’s utility function, alpha is the parameter used in the definition of KT’s utility, exp. refers to exponential. Shaded areas are interquantile (25-75%) margins and curves report the median values over 10 different runs.
I.8 Control on MuJoCo Environments

### v3

@@ 2 $\uparrow$ ? $\uparrow$
0 $\mathrel{\mkern-4.0mu\rightarrow\mkern-4.0mu}$ $\mathrel{\mkern-4.0mu\rightarrow\mkern-4.0mu}$ $\downarrow$
1 $\mathrel{\mkern-4.0mu\rightarrow\mkern-4.0mu}$ $\mathrel{\mkern-4.0mu\rightarrow\mkern-4.0mu}$
2 $\mathrel{\mkern-4.0mu\rightarrow\mkern-4.0mu}$ $\mathrel{\mkern-4.0mu\rightarrow\mkern-4.0mu}$ $\uparrow$
@@
0 $\mathrel{\mkern-4.0mu\rightarrow\mkern-4.0mu}$ $\mathrel{\mkern-4.0mu\rightarrow\mkern-4.0mu}$ ? $\downarrow$
1 $\uparrow$ ? $\mathrel{\mkern-4.0mu\rightarrow\mkern-4.0mu}$ $\downarrow$
2 $\uparrow$ $\uparrow$ $\mathrel{\mkern-4.0mu\rightarrow\mkern-4.0mu}$
3 $\uparrow$ $\uparrow$ ? $\mathrel{\mkern-4.0mu\rightarrow\mkern-4.0mu}$
@@
0 $\mathrel{\mkern-4.0mu\rightarrow\mkern-4.0mu}$ $\mathrel{\mkern-4.0mu\rightarrow\mkern-4.0mu}$ $\mathrel{\mkern-4.0mu\rightarrow\mkern-4.0mu}$ $\downarrow$
1 $\mathrel{\mkern-4.0mu\rightarrow\mkern-4.0mu}$ $\mathrel{\mkern-4.0mu\rightarrow\mkern-4.0mu}$ $\mathrel{\mkern-4.0mu\rightarrow\mkern-4.0mu}$ $\downarrow$
2 $\mathrel{\mkern-4.0mu\rightarrow\mkern-4.0mu}$ $\mathrel{\mkern-4.0mu\rightarrow\mkern-4.0mu}$ $\mathrel{\mkern-4.0mu\rightarrow\mkern-4.0mu}$
3 $\mathrel{\mkern-4.0mu\rightarrow\mkern-4.0mu}$ $\mathrel{\mkern-4.0mu\rightarrow\mkern-4.0mu}$ $\mathrel{\mkern-4.0mu\rightarrow\mkern-4.0mu}$ $\uparrow$
@@
Figure 23: Returns along the iterations of our PG algorithm for CPT-PO for traffic control. Shaded areas indicate a range of $\pm$ one standard deviation over 20 runs. See Appendix K for details.
Implementation details. In both cases, the risk-neutral optimal solution (going around the city center) is also a local optimum for the risk-averse objective, and, because it is a shorter path, is easier to stumble upon by chance when exploring the MDP. This means we have to implement special measures to force exploration. The algorithm used as is is prone to get stuck from time to time in local minima on this example. It would seem that our $w$ function, which is aggressively risk-averse, hinders exploration. To mitigate this, we introduce an entropy regularization term that we add to the score function with a decaying regularization weight in the policy gradient found in Theorem 3, see App. K for further details. We incorporate entropy regularization in the policy gradient as follows:
@@
K.7 Electricity Management
In this application involving continuous state and action spaces, we consider an individual home which has solar panels for producing electricity and a battery for storing energy (see Fig. 24, left). We consider a 24-hour time frame where the agent must decide the quantity of electricity to buy/sell, based on solar panel production, market prices, and battery levels. We use public data for selling prices recorded on a national electricity network. Public data is available online.555www.services-rte.com/en/view-data-published-by-rte/france-spot-electricity-exchange.html We experiment with risk-neutral, risk-averse, and risk-seeking objectives with 3 weight functions $w$ and a Gaussian neural network policy. Our algorithm performs well across these scenarios, with the risk-averse policy minimizing downside risk, and the risk-seeking policy maximizing potential gains. Fig. (24) (right) shows the distribution of total returns for different objectives. The most rewarding time to sell electricity is around 4pm (see prices in Fig. 24, center). However, selling too much too soon exposes us to the risk of falling short of battery during the night and risking to buy it later for a higher price. The risk-averse policy avoids selling a lot of electricity and tends to keep it stored until the end of the day. Conversely, the risk-seeking policy aggressively sells energy when the markets are high at the cost of possibly having to buy it again later in the day. The risk-averse policy has the distribution with the best left tail (worst cases are not too bad), the risk-seeking distribution has the best right tail (best cases are particularly good). The risk-neutral policy has the best mean value.
Figure 24: (Left) Electricity management environment: arrows refer to electricity flow. (Center) Electricity prices in a typical day, the blue line (right-hand side scale) records the electricity price on the European market, the shaded area (left-hand side scale) represents the total electricity production. (Right) Density of the empirical returns obtained by deploying different trained PG policies from different initializations, density computed using 10,000 runs for each curve.
K.8 Trading in Financial Markets
@@
  • Rewards: The rewards we consider are given by the log values of the ratio of the portfolio valuations at times $t$ and t1.t-1\,. Borrowing interest rates and trading fees are also considered in the computation. The reward function can also be easily modified in the environment thanks to the implementation of the Gym Trading Environment which builds on the standard Gym environments.
@@
Results. We refer the reader to Fig. 3. We make a few observations:
@@
  • Influence of the parameter $\alpha$ in KT’s utility: Observe that the exponent $\alpha$ in the utility distorts the function and shifts the returns significantly. Lower values of $\alpha$ lead to higher returns in this setting where the returns (as per the ratio definition of the reward) are smaller than 1. This parameter $\alpha$ provides a degree of freedom to model the behavior of the agent as per their perception of the returns. Different values of $\alpha$ modify the curvature of the utility function (w.r.t. the reference point which is $x_{0}=0$ here) which is concave for gains and convex for losses.
K.9 Control on MuJoCo Environments